# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

### v1

@@ Abstract
In this paper, we study generalized cryptographic semi-regular sequences, which are expected to generic in the space of homogeneous polynomial sequences on which the coordinate rings have Krull dimension one. We provide an upper-bound on the complexity of the Gröbner basis computation for such a sequence.
@@ 1 Introduction
In the theory of polynomial rings over fields, Gröbner bases introduced first by Buchberger [8] provide various methods for investigating important properties of an ideal and its associated algebraic variety. Here, a Gröbner basis is defined as a special kind of its generator set, relative to a monomial ordering. Buchberger also presented the first algorithm (Buchberger’s algorithm) for computing Gröbner bases. Lazard’s formulation in [30] and [31] relates the problem of computing Gröbner bases to the problem of reducing so-called Macaulay matrices. FGLM [19] and Gröbner walk [12] were propsed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven [37] can be also applied to such a conversion. Faugère’s proposed his innovative algorithms $F_{4}$ [17] and $F_{5}$ [18] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms. We refer to [16] for a survey on signature-based algorithms including $F_{5}$. Thanks to Gröbner basis conversion algorithms, it is important to estimate the complexity of computing a Gröbner basis with respect to a specific monomial ordering. In this paper, we consider only the case of graded monomial orderings.
@@
In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [29, Section 2.2] for rigorous definitions. If the ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied. In the case of inhomogeneous ideals, the solving degree can be greater than the maximal degree of the reduced Gröbner basis, whence estimating it directly is difficult in general, see [10] and [29] for details. On the other hand, it is well-known (cf. [27, Prop. 4.3.18]) that a Gröbner basis of an inhomogeneous ideal with respect to a graded monomial ordering is easily obtained from a Gröbner basis of the ideal of homogenized generators. Therefore, the homogeneous case is very important, and we focus on that case. In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ (resp. $R^{\prime}=K[x_{0},x_{1},\ldots,x_{n}]$) be the polynomial ring of $n$ (resp. $n+1$) variables over $K$. For a possibly inhomogeneous ideal $I$ of $R$ (resp. $R^{\prime}$), we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set of monomials in $R$ (resp. $R^{\prime}$), following [9, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [24, Definition 4]). If $I$ is generated by $F\subset R$, we may write $\mathrm{max.GB.deg}_{\prec}(F)$ in place of it.
@@
For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (cf. [2, Proposition 1]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication. (We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [35, Theorem 1.72].) With this fact, we have in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$. Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating the complexity of Gröbner basis computation in the homogeneous case. The case on which we focus on is the case where the Krull dimension $\mathrm{Krull.dim}(R/I)$ is equal to or less than one; see e.g., [24, Section 1] for a summary of existing results when $\mathrm{Krull.dim}R/I\geq 2$. In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known, and it is given by $\sum_{j=1}^{\ell}(d_{j}-1)+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell:=\max\{m,n\}$, where we set $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$. When $I$ is in quasi stable position (equivalently $R/\langle I,x_{n}\rangle$ is Artinian, see Lemma 2.1.8 below), we can apply Hashemi-Seiler’s inequality
$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(\langle I,x_{n}\rangle),\mathrm{hilb}(I)\}$ (1.0.1)
@@
for the graded reverse lexicographic ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$. Note that $\mathrm{hilb}(\langle I,x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(\langle I,x_{n}\rangle)$ since $R/\langle I,x_{n}\rangle$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$. For the definition of $d_{\rm reg}$ and $d_{\rm reg}$, see Remark 2.1.2 below.
On the other hand, in [29], when $F=\{f_{1},\ldots,f_{m}\}\subset R$ with $m\geq n$ is obtained by homogenizing an affine semi-regular sequence in $K[x_{1},\ldots,x_{n-1}]^{m}$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(\langle F\rangle_{R})$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$. Under an additional assumption that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (see Definition 2.2.7), we also proved in [29] that the right hand side of (1.0.1) is upper-bounded by the degree of regularity for a cryptographic semi-regular sequence of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n+1$ variables.
In this paper, we shall prove that, if $K$ is large enough, the inequality (1.0.1) always holds after a linear coordinate change. Then, we also estimate the complexity of computing a Gröbner basis for $I$, assuming that $I$ is generated by a generalized cryptographic semi-regular sequence. Such a sequence is expected to be generic (see Definition 2.3.1 for definition), analogously to Fröberg’s conjecture [20]. Our results are summarized into the following theorem and proposition:
Theorem 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$. Assume that $I$ has finitely many projective zeros over $\overline{K}$, and suppose that the cardinality of $K$ is greater than the number of projective zeros over $\overline{K}$ of $I$. Choose an arbitrary linear form $\ell$ in $R$ such that it does not vanish at any projective zero over $\overline{K}$ of $I$ (such an $\ell$ always exists by our assumption $\mathrm{Krull.dim}(R/I)\leq 1$). Then, the following (1) holds:
  1. (1)
    There exists a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ sending $\ell$ to $x_{n}$ such that
    ${\rm max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}$ (1.0.2)
    for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, where $I^{\sigma}$ is the image of $I$ by $\sigma$.
Moreover, if $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then following (2) and (3) also hold:
@@
    $\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.0.3)
@@
    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.0.4)
@@
    The complexity of computing a Gröbner basis of $I^{\sigma}$ with respect to $\prec$ is upper-bounded by
    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right),$ (1.0.5)
    where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
Proposition 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Let $I$ be a proper homogeneous ideal of $R$. Assume that the Krull dimension of $R/I$ is $\leq 1$ and $R/\langle I,x_{n}\rangle$ is Artinian. If $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.0.1) becomes an equality. Hence, the inequality (1.0.2) in Theorem 1 (2) also becomes an equality.
@@
Theorem 1 can be viewed as a generalization of [29, Proposition 1 and Corollary 1]. The bound $D_{\rm new}$ given in (1.0.4) coincides with Lazard’s bound when $m\leq n$, while it can be sharper than Lazard’s one when $m>n$. We also note that, the complexity bound (1.0.5) is better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [2, Proposition 1], [36, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [35, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf. [29, Section 4.3]).
@@
By Theorem 1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [29, Section 4.3]. Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [9, Theorem 7]. Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [27, Section 3.1.1] for definition. The complexity of reducing the Macaulay matrix can be easily estimated as $O(m\binom{n+D}{D}^{\omega})$, without assuming $D=O(n)$. Suppose the following three conditions:

### v2

# Generalization of semi-regular sequences:

@@ Abstract
Nowadays, the notion of semi-regular sequences, originally proposed by Fröberg, becomes very important not only in Mathematics, but also in Information Science, in particular Cryptology. For example, it is highly expected that randomly generated polynomials form a semi-regular sequence, and based on this observation, secure cryptosystems based on polynomial systems can be devised. In this paper, we deal with a semi-regular sequence and its extension, named a generalized cryptographic semi-regular sequence, and give precise analysis on the complexity of computing a Gröbner basis of the ideal generated by such a sequence with help of several regularities of the ideal related to Lazard’s bound on maximal Gröbner basis degree and other bounds. We also study the genericness of the property that a sequence is semi-regular, and its variants related to Fröberg’s conjecture. Moreover, we discuss on the genericness of another important property that the initial ideal is weakly reverse lexicographic, related to Moreno-Socías’ conjecture, and show some criteria to examine whether both Fröberg’s conjecture and Moreno-Socías’ one hold at the same time.
@@ 1 Introduction
In the theory of polynomial rings over fields, Gröbner bases introduced first by Buchberger [9] provide various methods for investigating important properties of an ideal and its associated algebraic variety. Here, a Gröbner basis is defined as a special kind of its generator set, relative to a monomial ordering. Buchberger also presented the first algorithm (Buchberger’s algorithm) for computing Gröbner bases. Lazard’s formulation in [38] and [39] relates the problem of computing Gröbner bases to the problem of reducing so-called Macaulay matrices. FGLM [22] and Gröbner walk [13] were proposed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven algorithm [46] can be also applied to such a conversion. Faugère proposed his innovative algorithms $F_{4}$ [20] and $F_{5}$ [21] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms. We refer to [19] for a survey on signature-based algorithms including $F_{5}$. Thanks to Gröbner basis conversion algorithms, it is important to estimate the complexity of computing a Gröbner basis with respect to a specific monomial ordering. (Throughout this paper, all the complexities are measured by the number of arithmetic operations in a field to which all the coefficients of the input polynomials belong.) In this paper, we consider only the case of graded monomial orderings; in particular, we adopt a graded reverse lexicographic ordering.
@@
In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [36, Section 2.2] for rigorous definitions. If the given ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied. In the case of inhomogeneous ideals, the solving degree can be greater than the maximal degree of the reduced Gröbner basis, whence estimating it directly is difficult in general, see [11] and [36] for details. On the other hand, it is well-known (cf. [34, Prop. 4.3.18]) that a Gröbner basis of an inhomogeneous ideal with respect to a graded monomial ordering is easily obtained from a Gröbner basis of the ideal of homogenized generators. Therefore, the homogeneous case is essential, and we focus on that case. In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$. The term “degree” means total degree, unless otherwise noted. Let $R_{d}$ be its homogeneous part of degree $d$ of $R$ (i.e., the $K$-linear space spanned by the monomials in $R$ of degree $d$) for each non-negative integer $d$. For an ideal (possibly inhomogeneous) $I$ of $R$, we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set $\mathcal{T}$ of monomials in $R$, following [10, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [31, Definition 4]). If $I$ is generated by $F\subset R$, we may write $\mathrm{max.GB.deg}_{\prec}(F)$ in place of it.
Lazard’s bound and Hashemi-Seiler’s inequality
@@
For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (see e.g., [3, Proposition 1], [45, Equation (24)]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication. (We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [44, Theorem 1.72].) From this fact, it follows in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$. Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating shapely the complexity of Gröbner basis computation in the homogeneous case. The case on which we focus is the case where the Krull dimension of $R/I$, denoted by $\mathrm{Krull.dim}(R/I)$, is equal to or less than one; see e.g., [31, Section 1] for a summary of existing results when $\mathrm{Krull.dim}(R/I)\geq 2$. In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known. Specifically, if $K$ is large enough, then there exists a (suitable) linear coordinate change $\sigma$ such that $R/(I^{\sigma}+\langle x_{n}\rangle)$ is Artinian (equivalently $I$ is in quasi stable position, see Lemma 2.1.8 below). From this, we assume without loss of generality that $R/(I+\langle x_{n}\rangle)$ is Artinian. In this case, putting $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$ and supposing that $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ (descending order), it follows that
$\mathrm{max.GB.deg}_{\prec}(I)\leq\sum_{j=1}^{\min\{m,n\}}(d_{j}-1)+1$ (1.1)
for the graded reverse lexicographical ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, see [39, Théorème 3.3] and [38, Theorem 2]. Recently, Hashemi-Seiler [31] re-proved Lazard’s bound by using the inequality
$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(I+\langle x_{n}\rangle),\mathrm{hilb}(I)\},$ (1.2)
@@
where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$. Note that $\mathrm{hilb}(I+\langle x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(I+\langle x_{n}\rangle)$ since $R/(I+\langle x_{n}\rangle)$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$. For the definition of $d_{\rm reg}$ and $\widetilde{d}_{\rm reg}$, see Remark 2.1.2 below.
Remark 1.
Note that the condition $\mathrm{Krull.dim}(R/I)\leq 1$ does not necessarily imply $d_{\rm reg}(I+\langle x_{n}\rangle)=\mathrm{hilb}(I+\langle x_{n}\rangle)<\infty$, without considering any linear coordinate change $\sigma$. For example, putting $F:=\{f_{1}:=x_{1}x_{2}-1,f_{2}:=x_{2}-1\}\subset K[x_{1},x_{2}]$ (inhomogeneous), it is obvious that $\mathrm{Krull.dim}(R[y]/I)=1$ for $I=\langle F^{h}\rangle$ but $d_{\rm reg}(I+\langle y\rangle)=d_{\rm reg}(\langle F^{\rm top}\rangle)=\infty$, where $F^{h}$ is the homogenization of $F$ by an extra variable $y$, and where $F^{\rm top}$ is the homogeneous part of highest degree, see Remark 3 below. In this case, $F^{h}=\{x_{1}x_{2}-y^{2},x_{2}-y\}$ and $F^{\rm top}=\{x_{1}x_{2},x_{2}\}$. On the other hand, if $K$ is a finite field of order $q$ and if we would compute a zero of $\langle F\rangle$ over $K$, we often add the set of so-called field equations $x_{i}^{q}-x_{i}$ to an inhomogeneous $F$, say $F_{1}:=F\cup\{x_{i}^{q}-x_{i}:1\leq i\leq n\}$, so that we always have $d_{\rm reg}(\langle F_{1}^{h}\rangle+\langle y\rangle)=d_{\rm reg}(\langle F_{1}^{\rm top}\rangle)<\infty$.
Our results and conjectures
In this paper, we focus on several cases that are expected to be generic in some sense, and obtain upper-bounds on $\mathrm{max.GB.deg}_{\prec}(I)$ and some results related to Fröberg’s conjecture [23] and Moreno-Socías’ one [41], in those cases. For example, we will assume that $R/(I+\langle x_{n}\rangle)$ has the Hilbert series associated to generic forms, as conjectured by Fröberg [23] (see Conjecture 2.3.7 below). Note that a homogeneous polynomial sequence is called cryptographic semi-regular (cf. [8]) if its associated coordinate ring has the Hilbert series conjectured by Fröberg, see Definition 2.2.3 and Proposition 2.2.4 below for details. In the literature, a cryptographic semi-regular sequence plays a very important role in analyzing Gröbner basis computation, both in theory and in practical applications to cryptography, see e.g., Bardet et al’s works [1], [2], [3], [4], Gorla et al.’s works [8], [10], [11], and our ones [35], [36]. In our earlier work [36], when $m\geq n$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(F)$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$ (ascending order), supposing that $(f_{1}|_{x_{n}=0},\ldots,f_{i}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ for each $1\leq i\leq m$. In [36], we also defined a generalized cryptographic semi-regular sequence (see Definition 2.2.8 below for the definition and Remark 2.2.9 for a construction), as an extension of cryptographic semi-regular sequence to the case of Krull dimension one.
As the first result of this paper, we prove Theorem 1.1, which provides an upper-bound on $\mathrm{max.GB.deg}_{\prec}(I)$ for a generalized cryptographic semi-regular sequence and a complexity bound on computing a Gröbner basis of $I$.
Theorem 1.1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and $\prec$ the graded reverse lexicographical ordering with $x_{n}\prec\cdots\prec x_{1}$. Let $I$ be a homogeneous ideal of $R$ generated by homogeneous polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus K$ such that $\mathrm{Krull.dim}(R/I)\leq 1$. Assume that $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ (equivalently $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular), and that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, Then the following (1) and (2) hold:
@@
    $\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.3)
@@
    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.4)
@@
    The complexity of computing a Gröbner basis of $I$ with respect to $\prec$ is upper-bounded by
    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right)$ (1.5)
    arithmetic operations in $K$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
The bound $D^{(n,m)}$ in (1.3) is optimal for a homogeneous ideal $I$ satisfying the assumptions of Theorem 1.1, see Example 3.1.4 below. The proof of (2) in Theorem 1.1 is based on a well-known fact that the a Gröbner basis of $I$ is obtained by computing the reduced row echelon form of the degree-$D$ Macaulay matrix for any $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(I)$, see Section 3.1 for details.
@@
Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf. [36, Section 4.3]). The bound $D^{(n,m)}$ given in (1.4) coincides with Lazard’s bound when $m\in\{n-1,n\}$, while it can be sharper than Lazard’s one when $m>n$. We also note that, the complexity bound (1.5) is slightly better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [3, Proposition 1], [45, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [44, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
As the second result, we shall prove that Hashemi-Seiler’s inequality (1.2) is as possible as tight under some assumptions, one of which is that an initial ideal is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well. Moreno-Socías [41] conjectured that, for the graded reverse lexicographical ordering, the initial ideal of a generic ideal is a weakly reverse lexicographic.
Theorem 1.2.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Let $I$ be a proper homogeneous ideal of $R$ with $\mathrm{Krull.dim}(R/I)\leq 1$ and $\mathrm{Krull.dim}(R/(I+\langle x_{n}\rangle))=0$. If $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ (equivalently $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular) and if the initial ideal ${\rm in}_{\prec}(I)=\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.2) becomes an equality.
After proving Theorems 1.1 and 1.2, in Section 4, we also study the genericness of our assumptions used in the theorems, based on the theory of parametric Gröbner bases. For this, in Section 2.3, we formulate the genericness of cryptographic semi-regular sequences in terms of parametric Gröbner bases. Then, in Section 4, some analogous results for generalized cryptographic semi-regular sequences will be proved, and the following conjectures will be raised as variants of Fröberg’s conjecture:
Conjecture 1.3 (Conjecture 4.1.10; see also [36, Conjecture 4.3.4]).
Assume that $K$ is an infinite field, and we endow $V_{n,m,(d_{1},\ldots,d_{m})}:=R_{d_{1}}\times\cdots\times R_{d_{m}}$ naturally with the Zariski topology. For any $n$, $m$, and $(d_{1},\ldots,d_{m})$ and for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $Z_{\bm{p}}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\bm{p}\in V_{\overline{K}}(\bm{H})\}$, namely there exists a non-empty Zariski-open set of $Z_{\bm{p}}$ with respect to the induced topology on which the property holds.
Conjecture 1.4 (Conjecture 4.1.11).
Assume that $K$ is an infinite field. For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $V_{n,m,(d_{1},\ldots,d_{m})}^{\dim\geq 1}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\mathrm{Krull.dim}(R/\langle\bm{H}\rangle)\geq 1\}$, namely there exists a non-empty Zariski-open set of $V_{n,m,(d_{1},\ldots,d_{m})}^{\dim\geq 1}$ with respect to the induced topology on which the property holds.
We also show a new criterion to examine whether Fröberg’s conjecture and Moreno-Socías’ one hold at the same time, for a fixed parameter set $(n,m,d_{1},\ldots,d_{m})$ (although these conjectures are set for any parameter set, see Remark 2 for details):
Theorem 1.5 (Theorem 4.2.1).
Assume that $K$ is an infinite field, and let $n,m,d_{1},\ldots,d_{m}$ be positive integers. If there exists a cryptographic semi-regular sequence $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ such that ${\rm in}_{\prec}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, then there also exists a non-empty Zariski-open set $U$ of $V_{n,m,(d_{1},\ldots,d_{m})}$ such that $\bm{H}$ is cryptographic semi-regular and ${\rm in}_{\prec}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}\in U$. Namely, Fröberg’s conjecture and Moreno-Socías’ one both hold for the parameter set $(n,m,d_{1},\ldots,d_{m})$.
Remark 2.
Pardue proved in his seminal paper [43] that Fröberg’s conjecture is equivalent to his several conjectures, and that Moreno-Socías’ conjecture implies all of them. Note that these equivalence and implication are intended to mean that each conjecture holds for any choice of parameters $(n,m,d_{1},\ldots,d_{m})$. More precisely, we see from the proof of his theorem [43, Theorem 2] that (1) and (2) are equivalent to each other and that (3) implies (1) and (2), over an infinite field:
  1. (1)
    (Fröberg’s conjecture) The property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is cryptographic semi-regular is generic (in terms of Definition 2.3.1) for any $(n,m,d_{1},\ldots,d_{m})$.
  2. (2)
    (Pardue’s conjecture [43, Conjecture B]) The property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is semi-regular is generic for any $(n,m,d_{1},\ldots,d_{m})$.
  3. (3)
    (Moreno-Socías’ conjecture [41]) The property that the initial ideal of the homogeneous ideal generated by $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ with respect to a graded reverse lexicographical ordering is weakly reverse lexicographic is generic for any $(n,m,d_{1},\ldots,d_{m})$.
Namely, even if the property in (3) is shown to be generic for a fixed parameter set $(n,m,d_{1},\ldots,d_{m})$, it is unclear that the property in (1) (or (2)) is generic for the same parameter set (this also applies to the case of (1) $\Rightarrow$ (2)).
Analogously to Theorem 1.5, we finally prove the following theorem, which can also examine Conjecture 1.3:
Theorem 1.6 (Proposition 4.2.3 and Proposition 4.2.4).
Under the same setting as in Theorem 1.5, assume that there exists a generalized cryptographic semi-regular sequence $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ with $\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)=1$ such that $\dim_{K}(R/\langle\bm{F}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)$ and ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic. Then, for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, there exists a non-empty Zariski-open set $U_{\bm{p}}$ of $Z_{\bm{p}}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\bm{p}\in V_{\overline{K}}(\bm{H})\}$ with respect to the induced topology such that $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}$ in $U_{\bm{p}}$.
To the end of Introduction, we remark an application of Theorem 1.1 to the inhomogeneous case.
@@
By Theorem 1.1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [36, Section 4.3]. Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [10, Theorem 7]. Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [34, Section 3.1.1] for definition. The complexity of reducing the Macaulay matrix can be easily estimated as $O(m\binom{n+D}{D}^{\omega})$, without assuming $D=O(n)$. Suppose the following three conditions: